# MapTracker: Tracking with Strided Memory Fusion for Consistent Vector HD Mapping

### v1

@@ Abstract
This paper presents a vector HD-mapping algorithm that formulates the mapping as a tracking task and uses a history of memory latents to ensure consistent reconstructions over time. Our method, MapTracker, accumulates a sensor stream into memory buffers of two latent representations: 1) Raster latents in the bird’s-eye-view (BEV) space and 2) Vector latents over the road elements (i.e., pedestrian-crossings, lane-dividers, and road-boundaries). The approach borrows the query propagation paradigm from the tracking literature that explicitly associates tracked road elements from the previous frame to the current, while fusing a subset of memory latents selected with distance strides to further enhance temporal consistency. A vector latent is decoded to reconstruct the geometry of a road element. The paper further makes benchmark contributions by 1) Improving processing code for existing datasets to produce consistent ground truth with temporal alignments and 2) Augmenting existing mAP metrics with consistency checks. MapTracker significantly outperforms existing methods on both nuScenes and Agroverse2 datasets by over 8% and 19% on the conventional and the new consistency-aware metrics, respectively. The code will be available on our project page: https://map-tracker.github.io.
@@ 5 Experiments
Table 1: Results on nuScenes[2]. The first column shows three different ground truth used for training and testing. “Consistent” is our temporarily consistent ground truth. The standard AP scores are reported for pedestrian crossing, lane-divider, road-boundary, and their average. C-mAP is our consistency-aware metric, which requires tracking information in the ground truth and is reported only for Consistent. ${}^{+}$: Numbers are from the original papers. ${}^{\dagger}$: Epochs for our multi-frame training.
@@
MapTR ${}^{+}$[18] R50 110 56.2 59.8 60.1 58.7 -
${}^{+}$ [7] SwinT 110 62.6 68.0 69.7 66.8
${}^{+}$[19] R50 110 68.1 68.3 69.7 68.7
${}^{+}$ [46] R50 110 67.1 69.8 71.4 69.4
@@
${}^{+}$ [36] R50 24 63.6 66.6 64.8 65.0
MapTracker (Ours) R50 ${}^{\dagger}$ 77.3 72.4 74.2 74.7
@@ 5.1 Quantitative evaluations
nuScenes results. Table 1 shows that both MapTRv2 and StreamMapNet achieve better mAP with our GT, which is expected as we fixed the inconsistencies in their original GT (explained in §4.1). StreamMapNet’s improvement is slightly higher since it has temporal modeling (whereas MapTR does not) and exploits temporal consistency in the data. MapTracker significantly outperforms the competing methods, especially with our consistent GT by more than 8% and 22% in the original and the consistency-aware mAP scores. Note that MapTracker is the only system to produce explicit tracking information (i.e., correspondences of reconstructed elements across frames), which is required for the consistency-area mAP. A simple matching algorithm creates tracks for the baseline methods (See Appendix for details).
Table 2: Results on Argoverse2 [37]. ${}^{+}$: Numbers are from the original papers. ${}^{\dagger}$: Epochs for our multi-frame training.
@@
MapTR ${}^{+}$[19] R50 6*4 62.9 72.1 67.1 67.4 -
${}^{+}$[46] R50 24*4 69.2 75.7 70.5 71.8
StmMapNet ${}^{+}$[41] R50 30 62.0 59.5 63.0 61.5 -
@@
${}^{+}$[36] R50 30 64.9 60.2 64.9 63.3
MapTracker (Ours) R50 ${}^{\dagger}$ 74.5 66.4 73.4 71.4
@@
Table 3: Results with geographically non-overlapping data proposed in StreamMapNet [41]. Our consistent ground truth is used. ${}^{\dagger}$: Epochs for our multi-frame training.
@@
nuScenes[2] StreamMapNet [41] R50 110 31.6 28.1 40.7 33.5 22.2
MapTracker (Ours) R50 ${}^{\dagger}$ 45.9 30.0 45.1 40.3 32.5
Argoverse2[37] StreamMapNet[41] R50 72 61.8 68.2 63.2 64.4 54.4
MapTracker (Ours) R50 ${}^{\dagger}$ 70.0 75.1 68.9 71.3 63.2
@@ 5.2 Results with geographically non-overlapping data
Official training/testing splits of nuScenes and Agroverse2 datasets have geographical overlaps (i.e., the same roads appear in both training/testing), which allows overfitting [20]. Table 3 compares the best baseline method StreamMapNet and MapTracker, based on geographically non-overlapping splits, proposed by StreamMapNet. MapTracker consistently outperforms with significant margins, demonstrating robust cross-scene generalization. Note that the performance for nuScenes datasets degrades for both methods. Upon careful inspection, the detection of road elements is successful but the regressed coordinates have large errors, leading to low performance. Appendix provides more analysis.
5.3 Ablation studies
Ablation studies in Table 4 demonstrate the contributions of key design elements in MapTracker. The first “baseline” entry is StreamMapNet without its temporal reasoning capabilities (i.e., without its BEV and vector streaming memories and modules). The second entry is StreamMapNet. Both methods are trained for 110 epochs till full convergence. The last three entries are the variants of MapTracker with or without the key design elements. The first variant drops the memory fusion components in the BEV/VEC modules. This variant utilizes the tracking formulation but relies on a single BEV/VEC memory to hold the past information. The second variant adds the memory buffers and the memory fusion components but without striding, that is, using the latest 4 frames for the fusion. This variant boosts performance, demonstrating the effectiveness of our memory mechanism. The last variant adds memory striding, which makes more effective use of the memory mechanism and improves performance.
@@ Appendix 0.D Additional Experimental Results and Analyses
Table 5: Full C-mAP on ${}^{\dagger}$: Epochs for our multi-frame training.
@@
MapTracker ${}^{\dagger}$varnothing (predicted) 75.4 65.0 66.9 69.1 72.5
@@ 0.D.1 Full C-mAP results
Table 6: Full C-mAP on Argoverse2[37]. ${}^{\dagger}$: Epochs for our multi-frame training.
@@
MapTracker ${}^{\dagger}$varnothing (predicted) 70.8 68.3 66.0 68.3 72.8

### v2

@@ Abstract
This paper presents a vector HD-mapping algorithm that formulates the mapping as a tracking task and uses a history of memory latents to ensure consistent reconstructions over time. Our method, MapTracker, accumulates a sensor stream into memory buffers of two latent representations: 1) Raster latents in the bird’s-eye-view (BEV) space and 2) Vector latents over the road elements (i.e., pedestrian-crossings, lane-dividers, and road-boundaries). The approach borrows the query propagation paradigm from the tracking literature that explicitly associates tracked road elements from the previous frame to the current, while fusing a subset of memory latents selected with distance strides to further enhance temporal consistency. A vector latent is decoded to reconstruct the geometry of a road element. The paper further makes benchmark contributions by 1) Improving processing code for existing datasets to produce consistent ground truth with temporal alignments and 2) Augmenting existing mAP metrics with consistency checks. MapTracker significantly outperforms existing methods on both nuScenes and Agroverse2 datasets by over 8% and 19% on the conventional and the new consistency-aware metrics, respectively. The code and models are available on our project page: https://map-tracker.github.io.
@@ 5 Experiments
Table 1: Results on nuScenes[2]. The first column shows three different ground truth used for training and testing. “Consistent” is our temporarily consistent ground truth. The standard AP scores are reported for pedestrian crossing, lane-divider, road-boundary, and their average. C-mAP is our consistency-aware metric, which requires tracking information in the ground truth and is reported only for Consistent. +: Numbers are from the original papers. : Epochs for our multi-frame training.
@@
MapTR MapTR+[18] R50 110 56.2 59.8 60.1 58.7 -
PivotNet+ [7] SwinT 110 62.6 68.0 69.7 66.8
MapTRv2+[19] R50 110 68.1 68.3 69.7 68.7
GeMap+ [46] R50 110 67.1 69.8 71.4 69.4
@@
SQD-MapNet+ [36] R50 24 63.6 66.6 64.8 65.0
MapTracker (Ours) R50 72 77.3 72.4 74.2 74.7
@@ 5.1 Quantitative evaluations
nuScenes results. Table 1 shows that both MapTRv2 and StreamMapNet achieve better mAP with our GT, which is expected as we fixed the inconsistencies in their original GT (explained in §4.1). StreamMapNet’s improvement is slightly higher since it has temporal modeling (whereas MapTR does not) and exploits temporal consistency in the data. MapTracker significantly outperforms the competing methods, especially with our consistent GT by more than 8% and 22% in the original and the consistency-aware mAP scores. Note that MapTracker is the only system to produce explicit tracking information (i.e., correspondences of reconstructed elements across frames), which is required for the consistency-area mAP. A simple matching algorithm creates tracks for the baseline methods (See Appendix for details). We also measure the running speed of
Table 2: Results on Argoverse2 [37]. +: Numbers are from the original papers. : Epochs for our multi-frame training.
@@
MapTR MapTRv2+[19] R50 6*4 62.9 72.1 67.1 67.4 -
GeMap+[46] R50 24*4 69.2 75.7 70.5 71.8
StmMapNet StreamMapNet+[41] R50 30 62.0 59.5 63.0 61.5 -
@@
SQD-MapNet+[36] R50 30 64.9 60.2 64.9 63.3
MapTracker (Ours) R50 35 74.5 66.4 73.4 71.4
@@
Table 3: Results with geographically non-overlapping data proposed in StreamMapNet [41]. Our consistent ground truth is used. : Epochs for our multi-frame training.
@@
60$\times$30m nuScenes[2] StreamMapNet [41] 110 31.6 28.1 40.7 33.5 22.2
MapTracker (Ours) 72 45.9 30.0 45.1 40.3 32.5
Argoverse2[37] StreamMapNet[41] 72 61.8 68.2 63.2 64.4 54.4
MapTracker (Ours) 35 70.0 75.1 68.9 71.3 63.2
100$\times$50m nuScenes[2] StreamMapNet [41] 110 25.1 18.9 25.0 23.0 14.6
MapTracker (Ours) 72 45.9 24.3 38.4 36.2 27.5
Argoverse2[37] StreamMapNet[41] 72 60.1 56.1 47.5 54.6 41.3
MapTracker (Ours) 35 71.2 64.6 58.5 64.8 55.7
@@ 5.2 Results with geographically non-overlapping data
Official training/testing splits of nuScenes and Agroverse2 datasets have geographical overlaps (i.e., the same roads appear in both training/testing), which allows overfitting [20]. Table 3 compares the best baseline method StreamMapNet [41] and MapTracker based on the geographically non-overlapping splits with two different perception ranges (the settings are proposed by StreamMapNet). MapTracker consistently outperforms StreamMapNet with significant margins, demonstrating robust cross-scene generalization. Note that the performance for nuScenes degrades for both methods. Upon careful inspection, road elements were detected successfully, but the regressed coordinates had large errors, leading to low performance.
5.3 More analysis
@@
Ablation study of core model components. Table 4 demonstrate the contributions of key design elements in MapTracker. The first “baseline” entry is StreamMapNet without its temporal reasoning capabilities (i.e., without its BEV and vector streaming memories and modules). The second entry is StreamMapNet. Both methods are trained for 110 epochs till full convergence. The last three entries are the variants of MapTracker with or without the key elements. The first variant drops the memory fusion components in the BEV/VEC modules. This variant utilizes the tracking formulation but relies on a single BEV/VEC memory to hold the past information, like the GRU embedding of StreamMapNet. The second variant adds the memory buffers and the memory fusion components but without the striding strategy, that is, using the latest 4 frames for the fusion. This variant significantly boosts performance, demonstrating the effectiveness of our memory mechanism. The last variant adds memory striding, which makes more effective use of the memory mechanism and improves performance.
Table 5: Ablation study of the four distance strides, evaluated on the old data split of the nuScenes dataset with 60m×$\times$30m range.
[…]
4 - 78.6 73.3 72.8 74.9 68.1
20 {0, 0, 0, 0} 78.8 73.5 72.9 75.0 68.2
{1, 2, 3, 4} 79.9 74.0 73.6 75.8 68.8
{1, 3, 5, 7} 79.9 74.0 73.8 75.9 68.8
{1, 5, 10, 15} 80.0 74.1 74.1 76.1 69.1
{5, 10, 15, 20} 77.6 71.3 72.6 73.8 51.5
Choice of memory strides. Table 5 presents the ablation study over the choice of the four distance strides. We believe {1m, 5m, 10m, 15m} works well for two main reasons: 1) The first entry (i.e., 1m) helps stabilize the prediction of the current frame and should be close to the current frame to align with the training setup (for each training sample, we randomly sample 4 previous frames from the last 10 frames to form a training clip); and 2) The entries should have proper distance gaps to minimize the information redundancy in the memory.
@@
Inference speed. We evaluate the inference speed of MapTRv2, StreamMapNet, and MapTracker on the nuScenes dataset using a single NVIDIA RTX 6000 GPU with batch size 1. The results are 12.5 FPS, 14.2 FPS, and 11.5 FPS, respectively. MapTracker has different query propagation from StreamMapNet and introduces additional memory fusion layers, so the FPS is lower than the original StreamMapNet by 19% (11.5 vs. 14.2). Our current implementation of the vector memory fusion layer uses for-loops to iterate through all vector instances without batching, which is sub-optimal and could be optimized for better running efficiency.
@@ 6 Conclusion
Acknowledgements. This research is partially supported by NSERC Discovery Grants, NSERC Alliance Grants, and John R. Evans Leaders Fund (JELF). We thank the Digital Research Alliance of Canada and BC DRI Group for providing computational resources.
@@ Appendix 0.D Additional Experimental Results and Analyses
Table 6: Full C-mAP on nuScenes[2].: Epochs for our multi-frame training.
@@
MapTracker 72 $\varnothing$ (predicted) 75.4 65.0 66.9 69.1 72.5
@@ 0.D.1 Full C-mAP results
Table 7: Full C-mAP on Argoverse2[37]. : Epochs for our multi-frame training.
@@
MapTracker 35 $\varnothing$ (predicted) 70.8 68.3 66.0 68.3 72.8